# Long-Tailed Out-of-Distribution Detection: Prioritizing Attention to Tail

### v2

@@ Abstract
Current out-of-distribution (OOD) detection methods typically assume balanced in-distribution (ID) data, while most real-world data follow a long-tailed distribution. Previous approaches to long-tailed OOD detection often involve balancing the ID data by reducing the semantics of head classes. However, this reduction can severely affect the classification accuracy of ID data. The main challenge of this task lies in the severe lack of features for tail classes, leading to confusion with OOD data. To tackle this issue, we introduce a novel Prioritizing Attention to Tail (PATT) method using augmentation instead of reduction. Our main intuition involves using a mixture of von Mises-Fisher (vMF) distributions to model the ID data and a temperature scaling module to boost the confidence of ID data. This enables us to generate infinite contrastive pairs, implicitly enhancing the semantics of ID classes while promoting differentiation between ID and OOD data. To further strengthen the detection of OOD data without compromising the classification performance of ID data, we propose feature calibration during the inference phase. By extracting an attention weight from the training set that prioritizes the tail classes and reduces the confidence in OOD data, we improve the OOD detection capability. Extensive experiments verified that our method outperforms the current state-of-the-art methods on various benchmarks. Code will be released soon.
@@ Introduction
Motivated by these observations, we propose a Prioritizing Attention to Tail (PATT) method for long-tailed OOD detection. The core idea of our design is: address the imbalance in ID data and subsequently enhance the distinction in confidence levels between ID and OOD samples. Specifically, we propose a temperature scaling-based implicit semantic augmentation contrastive learning (TISAC). Compared to traditional supervised contrastive learning (Khosla et al. 2020), TISAC incorporates two key insights tailored for long-tailed OOD detection: (1) implicit semantic augmentation contrastive learning effectively balances imbalanced ID data. Particularly, it models the ID data using a mixture of von Mises-Fisher (vMF) distributions, allowing us to sample a large number of contrastive pairs. Yet, sampling sufficient data from the vMF distributions at each training iteration is still inefficient, and we extend the number of samples to infinity and rigorously derive a closed-form for the expected contrastive loss. (2) temperature scaling, which ensures high confidence in ID data, thereby increasing the confidence gap between ID and OOD data. By doing so, our method improves accuracy for both head and tail classes as shown in Fig. 1 (a).
@@
Early LTR methods primarily involve re-sampling (Buda, Maki, and Mazurowski 2018; Byrd and Lipton 2019; He and Garcia 2009; Wallace et al. 2011) and re-weighting (Huang et al. 2016; Cui et al. 2019), which are effective but limited in addressing intra-class diversity in tail classes, leading to overfitting. To increase the intra-class diversity of tail data, many data augmentation methods have been employed, but explicit data augmentation methods (Bowles et al. 2018; Zhang et al. 2018; Yun et al. 2019) require significant time and resources. ISDA (Wang et al. 2019), an implicit data augmentation method that utilizes a mixture of Gaussian distribution to generate infinite samples, is an excellent solution to the aforementioned issue. Subsequently, RISDA (Chen et al. 2022) emerged, which effectively applies ISDA to long-tailed learning. Although these LTR methods exhibit excellent performance in the long-tailed classification, they lack specific designs for OOD samples.
@@
Samples from tail classes struggle to represent the diversity of tail class data. Additionally, as aforementioned, contrastive learning achieves classification by repelling instances from different classes and attracting same-class ones. Due to the severe imbalance in the training set, the repelling force from the head classes can cause the tail classes to be compressed together, making them difficult to separate. Naive re-weighting and re-sampling methods are difficult to achieve ideal results, as these approaches fail to address the semantic deficiency of tail classes. Inspired by implicit semantic data augmentation methods (Wang et al. 2019; Chen et al. 2022; Du et al. 2024), we propose a temperature scaling-based implicit semantic augmentation contrastive learning module (TISAC).
@@
During the inference phase, we element-wise multiply this weight $A^{scale}$ with feature $z$ to obtain the final calibrated feature $z^{cal}=z\odot A^{scale}$. By doing so, our model will prioritize attention to the features of tail classes and reduce the confidence of OOD data. Thus, feature calibration benefits both ID classification and OOD detection. Additionally, instead of using maximum softmax probability (MSP) (Hendrycks and Gimpel 2017) as OOD scores, we are inspired by energyOE (Liu et al. 2020) and use energy scores as OOD scores in the inference phase, which is defined as follows:
@@
We conduct experiments on widely used datasets, i.e., CIFAR10-LT, CIFAR100-LT (Cao et al. 2019), and ImageNet-LT (Liu et al. 2019) as ID training sets ($\mathcal{D}_{in}$). The standard CIFAR10, CIFAR100, and ImageNet test sets are used as ID test sets ($D^{test}_{in}$). Following PASCL (Wang et al. 2022), we utilize 300,000 samples from TinyImages80M (Torralba, Fergus, and Freeman 2008) as the surrogate OOD training data for CIFAR10/100-LT and ImageNet-Extra as the surrogate OOD training for ImageNet-LT. We set the default imbalance ratio to 100 for CIFAR10/100-LT. For OOD test data, we use Textures (Cimpoi et al. 2014), SVHN  (Netzer et al. 2011), Tiny ImageNet (Le and Yang 2015), LSUN (Yu et al. 2015), and Places365 (Zhou et al. 2017) introduced in the SC-OOD benchmark (Yang et al. 2021) as $\mathcal{D}_{test}^{out}$ for CIFAR10/100-LT. Additionally, we use CIFAR-100 as $\mathcal{D}_{test}^{out}$ for CIFAR10-LT and vice versa. We use ImageNet-1k-OOD as $\mathcal{D}_{test}^{out}$ for ImageNet-LT. More information about the datasets can be found in Appendix C.
@@ D.1 Additional Ablation Study
Table 11: Average results over six OOD test dataset on CIFAR100-LT with different temperature scaling hyperparameter $\varepsilon$

### v3

@@ Abstract
Current out-of-distribution (OOD) detection methods typically assume balanced in-distribution (ID) data, while most real-world data follow a long-tailed distribution. Previous approaches to long-tailed OOD detection often involve balancing the ID data by reducing the semantics of head classes. However, this reduction can severely affect the classification accuracy of ID data. The main challenge of this task lies in the severe lack of features for tail classes, leading to confusion with OOD data. To tackle this issue, we introduce a novel Prioritizing Attention to Tail (PATT) method using augmentation instead of reduction. Our main intuition involves using a mixture of von Mises-Fisher (vMF) distributions to model the ID data and a temperature scaling module to boost the confidence of ID data. This enables us to generate infinite contrastive pairs, implicitly enhancing the semantics of ID classes while promoting differentiation between ID and OOD data. To further strengthen the detection of OOD data without compromising the classification performance of ID data, we propose feature calibration during the inference phase. By extracting an attention weight from the training set that prioritizes the tail classes and reduces the confidence in OOD data, we improve the OOD detection capability. Extensive experiments verified that our method outperforms the current state-of-the-art methods on various benchmarks. Code is available at https://github.com/InaR-design/PATT.
@@ Introduction
Motivated by these observations, we propose a Prioritizing Attention to Tail (PATT) method for long-tailed OOD detection. The core idea of our design is: address the imbalance in ID data and subsequently enhance the distinction in confidence levels between ID and OOD samples. Specifically, we propose a temperature scaling-based implicit semantic augmentation contrastive learning (TISAC). Compared to traditional supervised contrastive learning (Khosla et al. 2020), TISAC incorporates two key insights tailored for long-tailed OOD detection: (1) implicit semantic augmentation contrastive learning effectively balances imbalanced ID data. Particularly, it models the ID data using a mixture of von Mises-Fisher (vMF) distributions, allowing us to sample a large number of contrastive pairs. Yet, sampling sufficient data from the vMF distributions at each training iteration is still inefficient, and we extend the number of samples to infinity and rigorously derive a closed-form for the expected contrastive loss. (2) temperature scaling, which ensures high confidence in ID data, thereby increasing the confidence gap between ID and OOD data. By doing so, our method improves accuracy for both head and tail classes as shown in Fig. 1 (a), where CE is a pure classification model, and the other three methods are designed for long-tailed OOD tasks.
@@
Early LTR methods primarily involve re-sampling (Buda, Maki, and Mazurowski 2018; Byrd and Lipton 2019; He and Garcia 2009; Wallace et al. 2011) and re-weighting (Huang et al. 2016; Cui et al. 2019), which are effective but limited in addressing intra-class diversity in tail classes, leading to overfitting. To increase the intra-class diversity of tail data, many data augmentation methods have been employed, but explicit data augmentation methods (Bowles et al. 2018; Zhang et al. 2018; Yun et al. 2019) require significant time and resources. ISDA (Wang et al. 2019), an implicit data augmentation method that utilizes a mixture of Gaussian distribution to generate infinite samples, is an excellent solution to the aforementioned issue. Subsequently, RISDA (Chen et al. 2022) and ProCo (Du et al. 2024)emerged, which effectively applies ISDA to long-tailed learning. Although these LTR methods exhibit excellent performance in the long-tailed classification, they lack specific designs for OOD samples.
@@
Samples from tail classes struggle to represent the diversity of tail data. Additionally, as aforementioned, contrastive learning achieves classification by repelling instances from different classes and attracting same-class ones. Due to the severe imbalance in the training set, the repelling force from the head classes can cause the tail classes to be compressed together, making them difficult to separate. Naive re-weighting and re-sampling methods are difficult to achieve ideal results, as these approaches fail to address the semantic deficiency of tail classes. Inspired by implicit semantic data augmentation methods (Wang et al. 2019; Chen et al. 2022; Du et al. 2024), we propose a temperature scaling-based implicit semantic augmentation contrastive learning module (TISAC), which consists of two components: Implicit Semantic Augmentation Contrastive Learning (ISAC) and Temperature Scaling-Based Logit Adjustment (TLA). ISAC is inspired by ISDA, which employs Gaussian distributions to model unconstrained features and obtains an upper bound on the expected loss. However, Our method diverges significantly from ISDA-based methods. Compared to Gaussian distribution, the vMF distribution focuses on feature direction rather than magnitude, which largely reduces model bias between classes. Additionally, the Gaussian distribution requires extensive data to compute covariance matrices, making it impractical to estimate Gaussian parameters for all classes with long-tailed data. In contrast, vMF’s mean direction vector and concentration parameter can be effectively calculated across batches. Moreover, we rigorously derived the expected loss instead of an upper bound. TLA is a temperature scaling-based version of logit adjustment, which, when combined with OE, creates a larger separation between ID and OOD data.
@@
During the inference phase, we element-wise multiply this weight $A^{scale}$ with feature $z$ to obtain the final calibrated feature $z^{cal}=z\odot A^{scale}$. By doing so, our model will prioritize attention to the features of tail classes and reduce the confidence of OOD data. Thus, feature calibration benefits both ID classification and OOD detection. Additionally, instead of using maximum softmax probability (MSP) as OOD scores, we are inspired by energyOE (Liu et al. 2020) and use energy scores as OOD scores in the inference phase, which is defined as follows:
@@
We conduct experiments on widely used datasets, i.e., CIFAR10-LT, CIFAR100-LT (Cao et al. 2019), and ImageNet-LT (Liu et al. 2019) as ID training sets ($\mathcal{D}_{in}$). The standard CIFAR10, CIFAR100, and ImageNet test sets are used as ID test sets ($D^{test}_{in}$). Following PASCL (Wang et al. 2022), we utilize 300,000 samples from TinyImages80M (Torralba, Fergus, and Freeman 2008) as the surrogate OOD training data for CIFAR10/100-LT and ImageNet-Extra as the surrogate OOD training for ImageNet-LT. We set the default imbalance ratio to 100 for CIFAR10/100-LT. For OOD test data, we use Textures (Cimpoi et al. 2014), SVHN  (Netzer et al. 2011), Tiny ImageNet (Le and Yang 2015), LSUN (Yu et al. 2015), and Places365 (Zhou et al. 2017) introduced in the SC-OOD benchmark (Yang et al. 2021) as $\mathcal{D}_{test}^{out}$ for CIFAR10/100-LT. Additionally, for near-OOD experiments (Fort, Ren, and Lakshminarayanan 2021), we use CIFAR-100 as $\mathcal{D}_{test}^{out}$ for CIFAR10-LT and vice versa. We use ImageNet-1k-OOD as $\mathcal{D}_{test}^{out}$ for ImageNet-LT. More information about the datasets can be found in Appendix C.
@@ Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (No. 62276221, No. 62376232); the Fujian Provincial Natural Science Foundation of China (No. 2022J01002).
@@ D.1 Additional Ablation Study
Table 11: Average results over six OOD test dataset on CIFAR100-LT with different temperature scaling hyperparameter $\varepsilon$.
IB AUROC AUPR-in AUPR-out FPR95 ACC
$[0,1]$ 89.66 89.96 88.43 30.81 84.80
$[0,2]$ 91.62 91.25 90.48 29.89 84.77
Table 12: Average results over six OOD test dataset on CIFAR10-LT with different interval boundary.
@@
Interval Boundary
We scale the attention weights to a range of [0,2] instead of [0,1] based on a simple idea that certain features should be amplified (falling within [1,2]) while others should be reduced (falling within [0,1]). Empirically, we conduct an experiment on CIFAR10-LT to compare the effectiveness of [0,2] and [0,1]. The results in the table below show a superiority of [0,2] over [0,1] across all OOD detection metrics, except that ID classification (ACC) is identical to [0,1]. These results verify that using [0,2] can yield better OOD detection performance.
@@ D.2 More Visualization
Method Time (s)
PASCL 2140
EAT 4248
PATT 2078
Table 14: Average Computation Trade-off over 5 replay on CIFAR10-LT with different method.
Computation Trade-off
Table 14 shows the total training time for each method on CIFAR10-LT using ResNet18 on an RTX-3090 GPU. Notably, PASCL and EAT are two-stage methods, whereas our PATT not only avoids additional training costs but also achieves end-to-end training.